### 24 Jan 2024 v3

# Possible high superconductivity in LaNiO under high pressure through manifestation of a nearly-half-filled bilayer Hubbard model

Abstract
Inspired by a recent experiment showing that LaNiO exhibits high superconductivity under high pressure, we theoretically revisit the possibility of superconductivity in this material. We find that superconductivity can take place which is somewhat similar to that of the bilayer Hubbard model consisting of the Ni orbitals. Although the coupling with the orbitals degrades superconductivity, can still be high enough to understand the experiment thanks to the very high reached in the bilayer Hubbard model.
pacs:
74.20.Mn,74.70.−bIntroduction.—Seeking for new unconventional high superconductors has been a great challenge ever since the discovery of the two families of unconventional high- superconductors, cuprates Keimer et al. (2015) and iron-based Hosono and Kuroki (2015). Several previous studies have shown that the cuprates are already in an ideal situation in that they are described by a single-orbital Hubbard model near half-filling on a square lattice, and hence their may be difficult to transcend Arita et al. (2000).
One possible approach for pursuing even higher is to realize in actual materials the bilayer Hubbard model, for which several studies have shown that the superconducting can be higher than that of the -wave superconducting state in the single-orbital Hubbard model Kuroki et al. (2002); Mishra et al. (2016). In fact, the bilayer Hubbard model has been widely studied from the past Bulut et al. (1992); Scalettar et al. (1994); Hetzel et al. (1994); dos Santos (1995); Liechtenstein et al. (1995); Kancharla and Okamoto (2007); Bouadim et al. (2008); Lanatà et al. (2008); Zhai et al. (2009); Maier and Scalapino (2011); Nakata et al. (2017); Maier et al. (2019); Matsumoto et al. (2020); Kato and Kuroki (2020); Kainth and Long (2019); Ma et al. (2022), and -wave superconductivity spm is found to be strongly enhanced near half-filling when the vertical electron hopping () between the layers is several times larger than the in-plane hopping, and the Fermi level () lies in the vicinity of the edge of one of the bands Bulut et al. (1992); Kuroki et al. (2002); Maier and Scalapino (2011); Mishra et al. (2016); Nakata et al. (2017); Maier et al. (2019); Matsumoto et al. (2020); Kato and Kuroki (2020). Nowadays, a band whose edge lies just below or above is often referred to as an incipient band, and has attracted interest in the study of iron-based superconductors Wang et al. (2011); Chen et al. (2015); Hirschfeld et al. (2011); Bang (2014, 2016, 2019); Charnukha et al. (2015); Miao et al. (2015), bilayer and ladder-type lattices Kuroki et al. (2005); Maier et al. (2019); Matsumoto et al. (2018); Ogura et al. (2017); Ogura (2019); Matsumoto et al. (2020); Kato and Kuroki (2020); Sakamoto and Kuroki (2020); Kainth and Long (2019), and flat band superconductivity Kobayashi et al. (2016); Misumi and Aoki (2017); Sayyad et al. (2020, 2023); Aoki (2020).
In fact, one of the present authors proposed that a double layer Ruddlesden-Popper compound LaNiO can be a good candidate for realizing the bilayer Hubbard model that satisfies the above-mentioned conditions Nakata et al. (2017). In this material, for which the Ni electron configuration is , the orbitals are elongated in the (out-of-plane) direction so that between the layers is much larger than the in-plane hoppings between the neighboring orbitals, and also the orbitals are nearly half-filled. Hence the portion of the electronic structure appears to be favorable for superconductivity from the above-mentioned viewpoint of the bilayer model, although deviation from the ideal model arises due to the presence of the Ni bands, which are nearly quarter-filled, overlapping and hybridizing with the bands.
Given this background, a recent experimental finding that LaNiO exhibits high superconductivity at high pressures Sun et al. (2023), which in itself has huge impact, is certainly intriguing. There, it was shown that the material undergoes a superconducting transition with a highest of 80 K under pressure above 14 GPa. Already several theoretical studies on this material, which have been performed independently from ours, have appeared right after the discovery of superconductivity Luo et al. (2023); Yang et al. (2023); Zhang et al. (2023); Lechermann et al. (2023). In particular, Ref. Yang et al. (2023) finds dominance of -wave pairing using functional renormalization group, while Ref. Lechermann et al. (2023) concludes -wave and -wave being dominant and subdominant, respectively.
Inspired by this experiment, here we theoretically revisit the possibility of superconductivity in LaNiO by constructing a four-orbital model that takes into account the crystal structure at high pressures. We find that -pairing superconductivity, which is somewhat similar to that of the bilayer Hubbard model, can take place with high that is consistent with the experimental observation. Although the coupling between the and the orbitals degrades superconductivity, can still be high because of the very high attained in the bilayer Hubbard model. We also discuss ways to further enhance superconductivity of this material.
Method.—First, we perform first-principles calculation to obtain the band structure of LaNiO using the QUANTUM ESPRESSO code Giannozzi et al. (2017). Perdew-Burke-Ernzerhof parametrization of the generalized gradient approximation (PBE-GGA) Perdew et al. (1996) and the scalar-relativistic version of the optimized norm-conserving Vanderbilt pseudopotentials Hamann (2013) taken from PseudoDojo van Setten et al. (2018) are used. As for the structural parameters, we adopt those given in Ref. Sun et al. (2023) for LaNiO under the pressure of GPa, namely, the experimentally determined lattice constants as well as the atomic positions determined by theoretical optimization. Since the orthorhombicity at GPa is quite small (, where are lattice constants for space group ), we adopt a body-centered tetragonal structure (, Fig. 1(a)) as in LaCuO, with the in-plane lattice constant determined as an average of the original ones, i.e., (see supplemental material SM for details). We take 100 Ry plane-wave cutoff energy, a 12 12 12 -mesh, and an energy width of 0.02 Ry for Gaussian smearing.
We then extract maximally localized Wannier functions Marzari and Vanderbilt (1997); Souza et al. (2001) using the RESPACK code res ; Nakamura et al. (2021, 2016, 2009, 2008); Nohara et al. (2009); Fujiwara et al. (2003), by which we also obtain the hopping parameters among the Wannier functions. We construct a four-orbital model consisting of the and the like Wannier orbitals centered at two Ni sites per unit cell. Important parameter values are given in Table 1. Figure 1(c) shows superposed band structures given by first-principles calculation and Wannier interpolation, where precise fitting around the Fermi level is achieved.
We explore the possibility of superconductivity for the obtained low-energy four-orbital model within the fluctuation-exchange (FLEX) approximation Bickers et al. (1989); Bickers and White (1991). As the interaction term of the Hamiltonian, we only take the on-site interactions, namely, intraorbital(interorbital) Coulomb interactions (), Hund’s coupling , and pair hopping . We assume the orbital rotational symmetry, namely, we take the same value of for the and the orbitals, and . Since typical values for cuprate are (where eV is a typical first-principles value Sakakibara et al. (2010, 2012, 2014, 2013) of the nearest neighbor hopping among the orbitals), we take eV. We also take , i.e., eV and eV. We calculate the self-energy induced by the spin-fluctuation formulated as shown in the literatures Lichtenstein and Katsnelson (1998); Yada and Kontani (2005); Mochizuki et al. (2005) in a self-consistent calculation. The explicit formulae of the irreducible, spin, and charge susceptibilities describing the fluctuations are shown in Eqs.(2)-(4) of Ref. Sakakibara et al. (2012). The real part of the self-energy at the lowest Matsubara frequency is subtracted in the same manner with Ref. Ikeda et al. (2010) to maintain the band structure around the Fermi level obtained by first-principles calculation.
The obtained Green’s function and the pairing interaction, mediated mainly by spin fluctuations, are plugged into the linearized Eliashberg equation. Since the eigenvalue of the linearized Eliashberg equation reaches unity at , we adopt it as a measure of superconductivity at a fixed temperature, eV. For convenience, we will call the eigenfunction (with the largest eigenvalue) of the linearized Eliashberg equation at the lowest Matsubara frequency (=) the “superconducting gap function”. We take a 16164 -point mesh and 2048 Matsubara frequencies for the FLEX calculation. Further details on the model construction and the FLEX calculation are provided in the supplemental materials SM .
| [eV] | |||||
Results and Discussions.—In Fig. 2(a), we show the eigenvalue of the linearized Eliashberg equation at eV as a function of the band filling , denoted as “original model”. corresponds to the stoichiometric composition of the actual material, and is varied assuming a rigid band. In Fig. 2(d), we show the superconducting gap function of the present model at in the band representation. It can be seen that the gap function is large at portions of the band where the orbital component is large, and the bonding and antibonding portions of the bands (see Fig. 2(b)) have opposite signs of the gap.
In Fig. 2(a), we also show a yellow shade presenting the range of the typical values of for the high cuprates obtained in the same way Sakakibara and Kotani (2019); Sakakibara et al. (2020). It can be seen that the present model, for or larger, exhibits large values comparable to those of the cuprates, which implies that the calculation results are consistent with the experimental observation of K. To elaborate on this point, we further perform FLEX calculation at adopting more realistic values for the two-body interaction parameters obtained by constrained random phase approximation (cRPA) Christiansson et al. (2023), and plot the eigenvalue in Fig. 2(a). The obtained value is close to the original value, reinforcing our conclusion. For comparison, we also show the eigenvalue for the infinite layer nickelate obtained in our previous study Sakakibara et al. (2020), which falls below the lower bound of the cuprate region and hence also consistent with the experiment Li et al. (2019).
| hybridization | ||||
| original model | Y | Y | 0 | 0 |
| (i) | N | Y | 0 | 0 |
| (ii) | Y | N | 0 | 0 |
| (iii) | N | N | 0 | 0 |
| (iv) | Y | Y | eV | 0 |
| (v) | Y | Y | 0 | eV |
To understand the origin of the large values, we study three other models in the same manner, namely, models in which the following couplings between the and the orbitals are eliminated: (i) the interorbital interactions , , , (ii) the hybridization, and (iii) both the interorbital interactions and the hybridization. Definition of the models, including those discussed later, is summarized in Table 2. The band structure without the hybridization is also presented in Fig. 2(c). In model (iii), the and the orbitals are completely decoupled, so that the superconducting state is equivalent to that of the bilayer Hubbard model consisting solely of the orbitals. It can be seen that both the hybridization and the interorbital interactions degrade superconductivity of the bilayer Hubbard model, but since of the bilayer model is significantly large, of the original model (full model with both the interorbital interactions and the hybridization included) is still large enough to explain the experimental observation. We will further elaborate on the suppression of superconductivity due to inter-orbital coupling in the supplemental materials SM .
The nature of the superconducting gap of the original model (Fig. 2(d)) can be more clearly understood by comparing it with that of model (ii) (the model in which the two orbitals are decoupled in one-body level) shown in Fig. 2(e). Here, the gap has opposite signs between the bonding and antibonding bands. It is an -wave superconducting gap in the wide sense of the term in that it changes sign between the two bands, but the antibonding band does not form a Fermi surface. We stress that even when one of the bands does not intersect the Fermi level, the spin fluctuations with finite energy arise as a pairing glue Nakata et al. (2017). Further details on this point are given in the supplemental material SM . The overall resemblance of the superconducting gaps in Figs. 2(d) and (e) further confirms our picture that the superconductivity in the present model is orbital driven.
In this context, it is also intriguing to give a look into the present system from a strong coupling viewpoint. Calculating within the second order perturbation, the interlayer exchange coupling between the orbitals gives eV for eV, which is quite large compared with, for example, the nearest neighbor superexchange coupling in the cuprates. is also much larger than the intralayer hopping between the neighboring orbitals. Such a large should lead to opening of a spin gap, and induce interlayer pairing superconductivity Dagotto et al. (1992), whose gap function changes its sign between bonding and antibonding bands in momentum space. This strong coupling picture is indeed consistent with the FLEX results for both the pure bilayer Hubbard model Nakata et al. (2017) and the present model.
In Refs. Yamazaki et al. (2020); Kitamine et al. (2020), some of the present authors studied cases where superconductivity emerges or is enhanced due to the interorbital interactions between the and other orbitals. The effect of the interorbital interactions in the present model is the opposite, namely, they degrade superconductivity. A large difference is that there is a bonding-antibonding splitting in the bands in the present bilayer system, which might be the reason why the effect of the interorbital interactions is the opposite. Further study on the origin of the difference between the single and bilayer systems is underway.
Finally, we discuss possible ways to further enhance superconductivity. The band filling dependence presented in Fig. 2 suggests that may be enhanced by doping electrons. In case it is difficult to dope electrons in the actual material, here we propose alternative ways for achieving a similar effect. We consider a model in which (iv) the level offset between the and the orbitals is increased by eV or (v) is increased by eV (see also Table 2). As depicted in Fig. 3(a), in both models, the band filling dependence of appears to be shifted toward the left (i.e., toward the smaller regime), so that larger values of are attained at , i.e., the stoichiometric band filling. From a material designing viewpoint, increasing and/or might be achieved by considering mixed anion materials.
The effect of increasing and/or can be understood by counting the number of electrons occupying the orbitals (namely, summing up the orbital weight assuming the non-interacting band structure) for each case. In Fig. 3(b), we plot against [], which is the average number of electrons per orbital. It can be seen that is mainly determined by [] within these three models, which once again supports the picture that the present superconductivity is orbital driven. Here, increasing and/or results in self-doping of electrons from the to the orbitals (see Fig. 1(b)). Superconductivity is enhanced as [] approaches unity, that is, as the orbital approaches half-filling, so that the electron correlation effects are enhanced, and at the same time, the Fermi level approaches both the bonding band top and the anti-bonding band bottom, thereby shifting the spin fluctuations toward the lower energy regime and making them more effective as a pairing glue.
Summary.—To summarize, we have studied the possibility of superconductivity in LaNiO taking into account the crystal structure under high pressure. The system can be considered as a bilayer Hubbard model of the orbitals coupled with the orbitals through interorbital interactions and hybridization. Although the interorbital couplings degrade superconductivity, the can still be high enough to explain the experimental observation, thanks to the very high reached in the bilayer Hubbard model. We have also discussed possible ways to enhance the superconductivity. Electron doping is likely to enhance superconductivity, but in case this is not feasible, increasing and/or are alternative ways of achieving a similar effect. This is because these modifications result in a self-doping of electrons from the to the orbitals. Studies on material designing along this line is underway.
Acknowledgements.
We are supported by JSPS KAKENHI Grant No. JP22K03512 (H. S.) and JP22K04907 (K. K.). The computing resource is supported by the supercomputer system HOKUSAI in RIKEN, and the supercomputer system (system-B) in the Institute for Solid State Physics, the University of Tokyo.References
- Keimer et al. (2015) B. Keimer, S. A. Kivelson, M. R. Norman, S. Uchida, and J. Zaanen, Nature 518, 179 (2015).
- Hosono and Kuroki (2015) H. Hosono and K. Kuroki, Physica C: Superconductivity and its Applications 514, 399 (2015), superconducting Materials: Conventional, Unconventional and Undetermined.
- Arita et al. (2000) R. Arita, K. Kuroki, and H. Aoki, Journal of the Physical Society of Japan 69, 1181 (2000).
- Kuroki et al. (2002) K. Kuroki, T. Kimura, and R. Arita, Phys. Rev. B 66, 184508 (2002).
- Mishra et al. (2016) V. Mishra, D. J. Scalapino, and T. A. Maier, Scientific Reports 6, 32078 (2016).
- Bulut et al. (1992) N. Bulut, D. J. Scalapino, and R. T. Scalettar, Phys. Rev. B 45, 5577 (1992).
- Scalettar et al. (1994) R. T. Scalettar, J. W. Cannon, D. J. Scalapino, and R. L. Sugar, Phys. Rev. B 50, 13419 (1994).
- Hetzel et al. (1994) R. E. Hetzel, W. von der Linden, and W. Hanke, Phys. Rev. B 50, 4159 (1994).
- dos Santos (1995) R. R. dos Santos, Phys. Rev. B 51, 15540 (1995).
- Liechtenstein et al. (1995) A. I. Liechtenstein, I. I. Mazin, and O. K. Andersen, Phys. Rev. Lett. 74, 2303 (1995).
- Kancharla and Okamoto (2007) S. S. Kancharla and S. Okamoto, Phys. Rev. B 75, 193103 (2007).
- Bouadim et al. (2008) K. Bouadim, G. G. Batrouni, F. Hébert, and R. T. Scalettar, Phys. Rev. B 77, 144527 (2008).
- Lanatà et al. (2008) N. Lanatà, P. Barone, and M. Fabrizio, Phys. Rev. B 78, 155127 (2008).
- Zhai et al. (2009) H. Zhai, F. Wang, and D.-H. Lee, Phys. Rev. B 80, 064517 (2009).
- Maier and Scalapino (2011) T. A. Maier and D. J. Scalapino, Phys. Rev. B 84, 180513(R) (2011).
- Nakata et al. (2017) M. Nakata, D. Ogura, H. Usui, and K. Kuroki, Phys. Rev. B 95, 214509 (2017).
- Maier et al. (2019) T. A. Maier, V. Mishra, G. Balduzzi, and D. J. Scalapino, Phys. Rev. B 99, 140504(R) (2019).
- Matsumoto et al. (2020) K. Matsumoto, D. Ogura, and K. Kuroki, Journal of the Physical Society of Japan 89, 044709 (2020).
- Kato and Kuroki (2020) D. Kato and K. Kuroki, Phys. Rev. Res. 2, 023156 (2020).
- Kainth and Long (2019) M. Kainth and M.-W. Long, (2019), arXiv:1904.07138.
- Ma et al. (2022) T. Ma, D. Wang, and C. Wu, Phys. Rev. B 106, 054510 (2022).
- (22) Here, “-wave” means that the sign of the superconducting gap function is the opposite between the hole-like bonding and the electron-like antibonding band of the bilayer model, as shown in the right side of Fig. 1 in Ref. Kitamine et al. (2020).
- Wang et al. (2011) F. Wang, F. Yang, M. Gao, Z.-Y. Lu, T. Xiang, and D.-H. Lee, Europhysics Letters 93, 57003 (2011).
- Chen et al. (2015) X. Chen, S. Maiti, A. Linscheid, and P. J. Hirschfeld, Phys. Rev. B 92, 224514 (2015).
- Hirschfeld et al. (2011) P. J. Hirschfeld, M. M. Korshunov, and I. I. Mazin, Reports on Progress in Physics 74, 124508 (2011).
- Bang (2014) Y. Bang, New Journal of Physics 16, 023029 (2014).
- Bang (2016) Y. Bang, New Journal of Physics 18, 113054 (2016).
- Bang (2019) Y. Bang, Scientific Reports 9, 3907 (2019).
- Charnukha et al. (2015) A. Charnukha, D. V. Evtushinsky, C. E. Matt, N. Xu, M. Shi, B. Büchner, N. D. Zhigadlo, B. Batlogg, and S. V. Borisenko, Scientific Reports 5, 18273 (2015).
- Miao et al. (2015) H. Miao, T. Qian, X. Shi, P. Richard, T. K. Kim, M. Hoesch, L. Y. Xing, X.-C. Wang, C.-Q. Jin, J.-P. Hu, and H. Ding, Nature Communications 6, 6056 (2015).
- Kuroki et al. (2005) K. Kuroki, T. Higashida, and R. Arita, Phys. Rev. B 72, 212509 (2005).
- Matsumoto et al. (2018) K. Matsumoto, D. Ogura, and K. Kuroki, Phys. Rev. B 97, 014516 (2018).
- Ogura et al. (2017) D. Ogura, H. Aoki, and K. Kuroki, Phys. Rev. B 96, 184513 (2017).
- Ogura (2019) D. Ogura, Theoretical Study of Electron Correlation Driven Superconductivity in Systems with Coexisting Wide and Narrow Bands, Doctor theis, Osaka Univ., Toyonaka, Osaka (2019).
- Sakamoto and Kuroki (2020) H. Sakamoto and K. Kuroki, Phys. Rev. Res. 2, 022055(R) (2020).
- Kobayashi et al. (2016) K. Kobayashi, M. Okumura, S. Yamada, M. Machida, and H. Aoki, Phys. Rev. B 94, 214501 (2016).
- Misumi and Aoki (2017) T. Misumi and H. Aoki, Phys. Rev. B 96, 155137 (2017).
- Sayyad et al. (2020) S. Sayyad, E. W. Huang, M. Kitatani, M.-S. Vaezi, Z. Nussinov, A. Vaezi, and H. Aoki, Phys. Rev. B 101, 014501 (2020).
- Sayyad et al. (2023) S. Sayyad, M. Kitatani, A. Vaezi, and H. Aoki, J. Phys. Condens. Matter 35, 245605 (2023).
- Aoki (2020) H. Aoki, Journal of Superconductivity and Novel Magnetism 33, 2341 (2020).
- Sun et al. (2023) H. Sun, M. Huo, X. Hu, J. Li, Y. Han, L. Tang, Z. Mao, P. Yang, B. Wang, J. Cheng, D.-X. Yao, G.-M. Zhang, and M. Wang, Nature 621, 493 (2023).
- Luo et al. (2023) Z. Luo, X. Hu, M. Wang, W. Wú, and D.-X. Yao, Phys. Rev. Lett. 131, 126001 (2023).
- Yang et al. (2023) Q.-G. Yang, D. Wang, and Q.-H. Wang, Phys. Rev. B 108, L140505 (2023).
- Zhang et al. (2023) Y. Zhang, L.-F. Lin, A. Moreo, and E. Dagotto, Phys. Rev. B 108, L180510 (2023).
- Lechermann et al. (2023) F. Lechermann, J. Gondolf, S. Bötzel, and I. M. Eremin, Phys. Rev. B 108, L201121 (2023).
- Giannozzi et al. (2017) P. Giannozzi, O. Andreussi, T. Brumme, O. Bunau, M. B. Nardelli, M. Calandra, R. Car, C. Cavazzoni, D. Ceresoli, M. Cococcioni, N. Colonna, I. Carnimeo, A. D. Corso, S. de Gironcoli, P. Delugas, R. A. DiStasio, A. Ferretti, A. Floris, G. Fratesi, G. Fugallo, R. Gebauer, U. Gerstmann, F. Giustino, T. Gorni, J. Jia, M. Kawamura, H.-Y. Ko, A. Kokalj, E. Küçükbenli, M. Lazzeri, M. Marsili, N. Marzari, F. Mauri, N. L. Nguyen, H.-V. Nguyen, A. O. de-la Roza, L. Paulatto, S. Poncé, D. Rocca, R. Sabatini, B. Santra, M. Schlipf, A. P. Seitsonen, A. Smogunov, I. Timrov, T. Thonhauser, P. Umari, N. Vast, X. Wu, and S. Baroni, Journal of Physics: Condensed Matter 29, 465901 (2017).
- Perdew et al. (1996) J. P. Perdew, K. Burke, and M. Ernzerhof, Phys. Rev. Lett. 77, 3865 (1996).
- Hamann (2013) D. R. Hamann, Phys. Rev. B 88, 085117 (2013).
- van Setten et al. (2018) M. van Setten, M. Giantomassi, E. Bousquet, M. Verstraete, D. Hamann, X. Gonze, and G.-M. Rignanese, Computer Physics Communications 226, 39 (2018).
- (50) See Supplemental Material at https://github.com/hskkbr/la3ni2o7paper for details, which includes Refs. Wang et al. (2023); Geisler et al. (2023); LaBollita et al. (2023); Qu et al. (2023a); Jiang et al. (2023); Qu et al. (2023b); Lu et al. (2023); Oh and Zhang (2023); Kakoi et al. (2023).
- Marzari and Vanderbilt (1997) N. Marzari and D. Vanderbilt, Phys. Rev. B 56, 12847 (1997).
- Souza et al. (2001) I. Souza, N. Marzari, and D. Vanderbilt, Phys. Rev. B 65, 035109 (2001).
- (53) RESPACK and its user guide can be downloaded from https://sites.google.com/view/kazuma7k6r.
- Nakamura et al. (2021) K. Nakamura, Y. Yoshimoto, Y. Nomura, T. Tadano, M. Kawamura, T. Kosugi, K. Yoshimi, T. Misawa, and Y. Motoyama, Computer Physics Communications 261, 107781 (2021).
- Nakamura et al. (2016) K. Nakamura, Y. Nohara, Y. Yosimoto, and Y. Nomura, Phys. Rev. B 93, 085124 (2016).
- Nakamura et al. (2009) K. Nakamura, Y. Yoshimoto, T. Kosugi, R. Arita, and M. Imada, Journal of the Physical Society of Japan 78, 083710 (2009).
- Nakamura et al. (2008) K. Nakamura, R. Arita, and M. Imada, Journal of the Physical Society of Japan 77, 093711 (2008).
- Nohara et al. (2009) Y. Nohara, S. Yamamoto, and T. Fujiwara, Phys. Rev. B 79, 195110 (2009).
- Fujiwara et al. (2003) T. Fujiwara, S. Yamamoto, and Y. Ishii, Journal of the Physical Society of Japan 72, 777 (2003).
- Bickers et al. (1989) N. E. Bickers, D. J. Scalapino, and S. R. White, Phys. Rev. Lett. 62, 961 (1989).
- Bickers and White (1991) N. E. Bickers and S. R. White, Phys. Rev. B 43, 8044 (1991).
- Sakakibara et al. (2010) H. Sakakibara, H. Usui, K. Kuroki, R. Arita, and H. Aoki, Phys. Rev. Lett. 105, 057003 (2010).
- Sakakibara et al. (2012) H. Sakakibara, H. Usui, K. Kuroki, R. Arita, and H. Aoki, Phys. Rev. B 85, 064501 (2012).
- Sakakibara et al. (2014) H. Sakakibara, K. Suzuki, H. Usui, S. Miyao, I. Maruyama, K. Kusakabe, R. Arita, H. Aoki, and K. Kuroki, Phys. Rev. B 89, 224505 (2014).
- Sakakibara et al. (2013) H. Sakakibara, K. Suzuki, H. Usui, K. Kuroki, R. Arita, D. J. Scalapino, and H. Aoki, Journal of Physics: Conference Series 454, 012021 (2013).
- Lichtenstein and Katsnelson (1998) A. I. Lichtenstein and M. I. Katsnelson, Phys. Rev. B 57, 6884 (1998).
- Yada and Kontani (2005) K. Yada and H. Kontani, J. Phys. Soc. Jpn. 74, 2161 (2005).
- Mochizuki et al. (2005) M. Mochizuki, Y. Yanase, and M. Ogata, Phys. Rev. Lett. 94, 147005 (2005).
- Ikeda et al. (2010) H. Ikeda, R. Arita, and J. Kuneš, Phys. Rev. B 81, 054502 (2010).
- Momma and Izumi (2011) K. Momma and F. Izumi, Journal of Applied Crystallography 44, 1272 (2011).
- Christiansson et al. (2023) V. Christiansson, F. Petocchi, and P. Werner, Phys. Rev. Lett. 131, 206501 (2023).
- Sakakibara and Kotani (2019) H. Sakakibara and T. Kotani, Phys. Rev. B 99, 195141 (2019).
- Sakakibara et al. (2020) H. Sakakibara, H. Usui, K. Suzuki, T. Kotani, H. Aoki, and K. Kuroki, Phys. Rev. Lett. 125, 077003 (2020).
- Li et al. (2019) D. Li, K. Lee, B. Y. Wang, M. Osada, S. Crossley, H. R. Lee, Y. Cui, Y. Hikita, and H. Y. Hwang, Nature 572, 624 (2019).
- Dagotto et al. (1992) E. Dagotto, J. Riera, and D. Scalapino, Phys. Rev. B 45, 5744 (1992).
- Yamazaki et al. (2020) K. Yamazaki, M. Ochi, D. Ogura, K. Kuroki, H. Eisaki, S. Uchida, and H. Aoki, Phys. Rev. Res. 2, 033356 (2020).
- Kitamine et al. (2020) N. Kitamine, M. Ochi, and K. Kuroki, Phys. Rev. Res. 2, 042032(R) (2020).
- Wang et al. (2023) L. Wang, Y. Li, S. Xie, F. Liu, H. Sun, C. Huang, Y. Gao, T. Nakagawa, B. Fu, D. Bo, Z. Cao, R. Yu, S. I. Kawaguchi, H. Kobayashi, M. Wang, C. Jin, H. Mao, and H. Liu, (2023), arXiv:2311.09186.
- Geisler et al. (2023) B. Geisler, J. J. Hamlin, G. R. Stewart, R. G. Henning, and P. J. Hirshfeld, (2023), arXiv:2309.15078.
- LaBollita et al. (2023) H. LaBollita, V. Pardo, M. R. Norman, and A. S. Botana, (2023), arXiv:2309.17279.
- Qu et al. (2023a) X.-Z. Qu, D.-W. Qu, W. Li, and G. Su, (2023a), arXiv:2311.12769.
- Jiang et al. (2023) K. Jiang, Z. Wang, and F.-C. Zhang, (2023), arXiv:2308.06771.
- Qu et al. (2023b) X.-Z. Qu, D.-W. Qu, J. Chen, C. Wu, F. Yang, W. Li, and G. Su, (2023b), arXiv:2307.16873.
- Lu et al. (2023) C. Lu, Z. Pan, F. Yang, and C. Wu, (2023), arXiv:2307.14965.
- Oh and Zhang (2023) H. Oh and Y.-H. Zhang, Phys. Rev. B 108, 174511 (2023).
- Kakoi et al. (2023) M. Kakoi, T. Kaneko, H. Sakakibara, M. Ochi, and K. Kuroki, (2023), arXiv:2312.04304.

### 06 Mar 2024 v5

@@
 (2017).
-Perdew-Burke-Ernzerhof parametrization of the generalized gradient approximation (PBE-GGA) Perdew et al.
+The Perdew-Burke-Ernzerhof parametrization of the generalized gradient approximation (PBE-GGA) Perdew et al.
 (1996) and the scalar-relativistic version of the optimized norm-conserving Vanderbilt pseudopotentials Hamann (2013) taken from PseudoDojo van Setten et al.
 (2018) are used.
-As for the structural parameters, we adopt those given in Ref.
+As for the structural parameters, we adopt those given in the extended data table 1 of Ref.
 Sun et al.
@@
 Since the orthorhombicity at GPa is quite small (, where are lattice constants for space group ), we adopt a body-centered tetragonal structure (, Fig.
-1(a)) as in LaCuO, with the in-plane lattice constant determined as an average of the original ones, i.e., (see supplemental material SM for details).
+1(a)) as in LaCuO, with the in-plane lattice constant determined as an average of the original ones, i.e., (see supplemental material SM for details, which includes Refs.
+Qu et al.
+(2023); Jiang et al.
+(2024); Qu et al.
+(2024); Lu et al.
+(2023); Oh and Zhang (2023); Kakoi et al.
+(2023)).
+Note that if optimizing not only the internal coordinates but also the lattice parameters within PBE-GGA, the structure indeed goes to Geisler et al.
+(2023); LaBollita et al.
+(2023), in agreement with a recent experiment Wang et al.
+(2023).
 We take 100 Ry plane-wave cutoff energy, a 12 12 12 -mesh, and an energy width of 0.02 Ry for Gaussian smearing.
@@
 2(d), we show the superconducting gap function of the present model at in the band representation.
-It can be seen that the gap function is large at portions of the band where the orbital component is large, and the bonding and antibonding portions of the bands (see Fig.
-2(b)) have opposite signs of the gap.
+It can be seen that the gap function is large at portions of the band where the orbital component is large, and the bonding and antibonding portions of the bands [see Fig.
+2(b)] have opposite signs of the gap.
 In Fig.
@@
 K.).
-The computing resource is supported by the supercomputer system HOKUSAI in RIKEN, and the supercomputer system (system-B) in the Institute for Solid State Physics, the University of Tokyo.References
-- Keimer et al. (2015) B. Keimer, S. A. Kivelson, M. R. Norman, S. Uchida, and J. Zaanen, Nature 518, 179 (2015).
-- Hosono and Kuroki (2015) H. Hosono and K. Kuroki, Physica C: Superconductivity and its Applications 514, 399 (2015), superconducting Materials: Conventional, Unconventional and Undetermined.
-- Arita et al. (2000) R. Arita, K. Kuroki, and H. Aoki, Journal of the Physical Society of Japan 69, 1181 (2000).
+The computing resource is supported by the supercomputer system HOKUSAI in RIKEN, the supercomputer system (system-B) in the Institute for Solid State Physics, the University of Tokyo, and the supercomputer of Academic Center for Computing and Media Studies (ACCMS), Kyoto University.References
+- Keimer et al. (2015) B. Keimer, S. A. Kivelson, M. R. Norman, S. Uchida, and J. Zaanen, Nature(London) 518, 179 (2015).
+- Hosono and Kuroki (2015) H. Hosono and K. Kuroki, Physica(Amsterdam) 514C, 399 (2015), superconducting Materials: Conventional, Unconventional and Undetermined.
+- Arita et al. (2000) R. Arita, K. Kuroki, and H. Aoki, J. Phys. Soc. Jpn. 69, 1181 (2000).
 - Kuroki et al. (2002) K. Kuroki, T. Kimura, and R. Arita, Phys. Rev. B 66, 184508 (2002).
-- Mishra et al. (2016) V. Mishra, D. J. Scalapino, and T. A. Maier, Scientific Reports 6, 32078 (2016).
+- Mishra et al. (2016) V. Mishra, D. J. Scalapino, and T. A. Maier, Sci. Rep. 6, 32078 (2016).
 - Bulut et al. (1992) N. Bulut, D. J. Scalapino, and R. T. Scalettar, Phys. Rev. B 45, 5577 (1992).
@@
 - Maier et al. (2019) T. A. Maier, V. Mishra, G. Balduzzi, and D. J. Scalapino, Phys. Rev. B 99, 140504(R) (2019).
-- Matsumoto et al. (2020) K. Matsumoto, D. Ogura, and K. Kuroki, Journal of the Physical Society of Japan 89, 044709 (2020).
+- Matsumoto et al. (2020) K. Matsumoto, D. Ogura, and K. Kuroki, J. Phys. Soc. Jpn. 89, 044709 (2020).
 - Kato and Kuroki (2020) D. Kato and K. Kuroki, Phys. Rev. Res. 2, 023156 (2020).
@@
 (2020).
-- Wang et al. (2011) F. Wang, F. Yang, M. Gao, Z.-Y. Lu, T. Xiang, and D.-H. Lee, Europhysics Letters 93, 57003 (2011).
+- Wang et al. (2011) F. Wang, F. Yang, M. Gao, Z.-Y. Lu, T. Xiang, and D.-H. Lee, Europhys. Lett. 93, 57003 (2011).
 - Chen et al. (2015) X. Chen, S. Maiti, A. Linscheid, and P. J. Hirschfeld, Phys. Rev. B 92, 224514 (2015).
-- Hirschfeld et al. (2011) P. J. Hirschfeld, M. M. Korshunov, and I. I. Mazin, Reports on Progress in Physics 74, 124508 (2011).
+- Hirschfeld et al. (2011) P. J. Hirschfeld, M. M. Korshunov, and I. I. Mazin, Rep. Prog. Phys. 74, 124508 (2011).
 - Bang (2014) Y. Bang, New Journal of Physics 16, 023029 (2014).
 - Bang (2016) Y. Bang, New Journal of Physics 18, 113054 (2016).
-- Bang (2019) Y. Bang, Scientific Reports 9, 3907 (2019).
-- Charnukha et al. (2015) A. Charnukha, D. V. Evtushinsky, C. E. Matt, N. Xu, M. Shi, B. Büchner, N. D. Zhigadlo, B. Batlogg, and S. V. Borisenko, Scientific Reports 5, 18273 (2015).
-- Miao et al. (2015) H. Miao, T. Qian, X. Shi, P. Richard, T. K. Kim, M. Hoesch, L. Y. Xing, X.-C. Wang, C.-Q. Jin, J.-P. Hu, and H. Ding, Nature Communications 6, 6056 (2015).
+- Bang (2019) Y. Bang, Sci. Rep. 9, 3907 (2019).
+- Charnukha et al. (2015) A. Charnukha, D. V. Evtushinsky, C. E. Matt, N. Xu, M. Shi, B. Büchner, N. D. Zhigadlo, B. Batlogg, and S. V. Borisenko, Sci. Rep. 5, 18273 (2015).
+- Miao et al. (2015) H. Miao, T. Qian, X. Shi, P. Richard, T. K. Kim, M. Hoesch, L. Y. Xing, X.-C. Wang, C.-Q. Jin, J.-P. Hu, and H. Ding, Nat. Commun. 6, 6056 (2015).
 - Kuroki et al. (2005) K. Kuroki, T. Higashida, and R. Arita, Phys. Rev. B 72, 212509 (2005).
@@
 - Ogura et al. (2017) D. Ogura, H. Aoki, and K. Kuroki, Phys. Rev. B 96, 184513 (2017).
-- Ogura (2019) D. Ogura, Theoretical Study of Electron Correlation Driven Superconductivity in Systems with Coexisting Wide and Narrow Bands, Doctor theis, Osaka Univ., Toyonaka, Osaka (2019).
+- Ogura (2019) D. Ogura, Theoretical Study of Electron Correlation Driven Superconductivity in Systems with Coexisting Wide and Narrow Bands, Doctoral theis, Osaka University, Toyonaka, Osaka (2019).
 - Sakamoto and Kuroki (2020) H. Sakamoto and K. Kuroki, Phys. Rev. Res. 2, 022055(R) (2020).
@@
 - Sayyad et al. (2023) S. Sayyad, M. Kitatani, A. Vaezi, and H. Aoki, J. Phys. Condens. Matter 35, 245605 (2023).
-- Aoki (2020) H. Aoki, Journal of Superconductivity and Novel Magnetism 33, 2341 (2020).
-- Sun et al. (2023) H. Sun, M. Huo, X. Hu, J. Li, Y. Han, L. Tang, Z. Mao, P. Yang, B. Wang, J. Cheng, D.-X. Yao, G.-M. Zhang, and M. Wang, Nature 621, 493 (2023).
+- Aoki (2020) H. Aoki, J. Supercond. Novel Magn. 33, 2341 (2020).
+- Sun et al. (2023) H. Sun, M. Huo, X. Hu, J. Li, Y. Han, L. Tang, Z. Mao, P. Yang, B. Wang, J. Cheng, D.-X. Yao, G.-M. Zhang, and M. Wang, Nature(London) 621, 493 (2023).
 - Luo et al. (2023) Z. Luo, X. Hu, M. Wang, W. Wú, and D.-X. Yao, Phys. Rev. Lett. 131, 126001 (2023).
@@
 - Hamann (2013) D. R. Hamann, Phys. Rev. B 88, 085117 (2013).
-- van Setten et al. (2018) M. van Setten, M. Giantomassi, E. Bousquet, M. Verstraete, D. Hamann, X. Gonze, and G.-M. Rignanese, Computer Physics Communications 226, 39 (2018).
+- van Setten et al. (2018) M. van Setten, M. Giantomassi, E. Bousquet, M. Verstraete, D. Hamann, X. Gonze, and G.-M. Rignanese, Comput. Phys. Commun. 226, 39 (2018).
+- Momma and Izumi (2011) K. Momma and F. Izumi, J. Appl. Crystallogr. 44, 1272 (2011).
 - (51) See Supplemental Material at https://github.com/hskkbr/la3ni2o7paper for details, which includes Refs.
@@
 (2023); Qu et al.
-(2023a); Jiang et al.
+(2023); Jiang et al.
 (2024); Qu et al.
-(2023b); Lu et al.
+(2024); Lu et al.
 (2023); Oh and Zhang (2023); Kakoi et al.
 (2023).
-- Marzari and Vanderbilt (1997) N. Marzari and D. Vanderbilt, Phys. Rev. B 56, 12847 (1997).
-- Souza et al. (2001) I. Souza, N. Marzari, and D. Vanderbilt, Phys. Rev. B 65, 035109 (2001).
-- (53) RESPACK and its user guide can be downloaded from https://sites.google.com/view/kazuma7k6r.
-- Nakamura et al. (2021) K. Nakamura, Y. Yoshimoto, Y. Nomura, T. Tadano, M. Kawamura, T. Kosugi, K. Yoshimi, T. Misawa, and Y. Motoyama, Computer Physics Communications 261, 107781 (2021).
-- Nakamura et al. (2016) K. Nakamura, Y. Nohara, Y. Yosimoto, and Y. Nomura, Phys. Rev. B 93, 085124 (2016).
-- Nakamura et al. (2009) K. Nakamura, Y. Yoshimoto, T. Kosugi, R. Arita, and M. Imada, Journal of the Physical Society of Japan 78, 083710 (2009).
-- Nakamura et al. (2008) K. Nakamura, R. Arita, and M. Imada, Journal of the Physical Society of Japan 77, 093711 (2008).
-- Nohara et al. (2009) Y. Nohara, S. Yamamoto, and T. Fujiwara, Phys. Rev. B 79, 195110 (2009).
-- Fujiwara et al. (2003) T. Fujiwara, S. Yamamoto, and Y. Ishii, Journal of the Physical Society of Japan 72, 777 (2003).
-- Bickers et al. (1989) N. E. Bickers, D. J. Scalapino, and S. R. White, Phys. Rev. Lett. 62, 961 (1989).
-- Bickers and White (1991) N. E. Bickers and S. R. White, Phys. Rev. B 43, 8044 (1991).
-- Sakakibara et al. (2010) H. Sakakibara, H. Usui, K. Kuroki, R. Arita, and H. Aoki, Phys. Rev. Lett. 105, 057003 (2010).
-- Sakakibara et al. (2012) H. Sakakibara, H. Usui, K. Kuroki, R. Arita, and H. Aoki, Phys. Rev. B 85, 064501 (2012).
-- Sakakibara et al. (2014) H. Sakakibara, K. Suzuki, H. Usui, S. Miyao, I. Maruyama, K. Kusakabe, R. Arita, H. Aoki, and K. Kuroki, Phys. Rev. B 89, 224505 (2014).
-- Sakakibara et al. (2013) H. Sakakibara, K. Suzuki, H. Usui, K. Kuroki, R. Arita, D. J. Scalapino, and H. Aoki, Journal of Physics: Conference Series 454, 012021 (2013).
-- Lichtenstein and Katsnelson (1998) A. I. Lichtenstein and M. I. Katsnelson, Phys. Rev. B 57, 6884 (1998).
-- Yada and Kontani (2005) K. Yada and H. Kontani, J. Phys. Soc. Jpn. 74, 2161 (2005).
-- Mochizuki et al. (2005) M. Mochizuki, Y. Yanase, and M. Ogata, Phys. Rev. Lett. 94, 147005 (2005).
-- Ikeda et al. (2010) H. Ikeda, R. Arita, and J. Kuneš, Phys. Rev. B 81, 054502 (2010).
-- Momma and Izumi (2011) K. Momma and F. Izumi, Journal of Applied Crystallography 44, 1272 (2011).
-- Christiansson et al. (2023) V. Christiansson, F. Petocchi, and P. Werner, Phys. Rev. Lett. 131, 206501 (2023).
-- Sakakibara and Kotani (2019) H. Sakakibara and T. Kotani, Phys. Rev. B 99, 195141 (2019).
-- Sakakibara et al. (2020) H. Sakakibara, H. Usui, K. Suzuki, T. Kotani, H. Aoki, and K. Kuroki, Phys. Rev. Lett. 125, 077003 (2020).
-- Li et al. (2019) D. Li, K. Lee, B. Y. Wang, M. Osada, S. Crossley, H. R. Lee, Y. Cui, Y. Hikita, and H. Y. Hwang, Nature 572, 624 (2019).
-- Dagotto et al. (1992) E. Dagotto, J. Riera, and D. Scalapino, Phys. Rev. B 45, 5744 (1992).
-- Yamazaki et al. (2020) K. Yamazaki, M. Ochi, D. Ogura, K. Kuroki, H. Eisaki, S. Uchida, and H. Aoki, Phys. Rev. Res. 2, 033356 (2020).
-- Kitamine et al. (2020) N. Kitamine, M. Ochi, and K. Kuroki, Phys. Rev. Res. 2, 042032(R) (2020).
+- Qu et al. (2023) X.-Z. Qu, D.-W. Qu, W. Li, and G. Su, (2023), arXiv:2311.12769.
+- Jiang et al. (2024) K. Jiang, Z. Wang, and F.-C. Zhang, Chin. Phys. Lett. 41, 017402 (2024).
+- Qu et al. (2024) X.-Z. Qu, D.-W. Qu, J. Chen, C. Wu, F. Yang, W. Li, and G. Su, Phys. Rev. Lett. 132, 036502 (2024).
+- Lu et al. (2023) C. Lu, Z. Pan, F. Yang, and C. Wu, (2023), arXiv:2307.14965.
+- Oh and Zhang (2023) H. Oh and Y.-H. Zhang, Phys. Rev. B 108, 174511 (2023).
+- Kakoi et al. (2023) M. Kakoi, T. Kaneko, H. Sakakibara, M. Ochi, and K. Kuroki, (2023), arXiv:2312.04304.
+- Geisler et al. (2023) B. Geisler, J. J. Hamlin, G. R. Stewart, R. G. Henning, and P. J. Hirshfeld, (2023), arXiv:2309.15078.
+- LaBollita et al. (2023) H. LaBollita, V. Pardo, M. R. Norman, and A. S. Botana, (2023), arXiv:2309.17279.
 - Wang et al.
@@
 Liu, (2023), arXiv:2311.09186.
-- Geisler et al. (2023) B. Geisler, J. J. Hamlin, G. R. Stewart, R. G. Henning, and P. J. Hirshfeld, (2023), arXiv:2309.15078.
-- LaBollita et al. (2023) H. LaBollita, V. Pardo, M. R. Norman, and A. S. Botana, (2023), arXiv:2309.17279.
-- Qu et al. (2023a) X.-Z. Qu, D.-W. Qu, W. Li, and G. Su, (2023a), arXiv:2311.12769.
-- Jiang et al. (2023) K. Jiang, Z. Wang, and F.-C. Zhang, (2023), arXiv:2308.06771.
-- Qu et al. (2023b) X.-Z. Qu, D.-W. Qu, J. Chen, C. Wu, F. Yang, W. Li, and G. Su, (2023b), arXiv:2307.16873.
-- Lu et al. (2023) C. Lu, Z. Pan, F. Yang, and C. Wu, (2023), arXiv:2307.14965.
-- Oh and Zhang (2023) H. Oh and Y.-H. Zhang, Phys. Rev. B 108, 174511 (2023).
-- Kakoi et al. (2023) M. Kakoi, T. Kaneko, H. Sakakibara, M. Ochi, and K. Kuroki, (2023), arXiv:2312.04304.
+- Marzari and Vanderbilt (1997) N. Marzari and D. Vanderbilt, Phys. Rev. B 56, 12847 (1997).
+- Souza et al. (2001) I. Souza, N. Marzari, and D. Vanderbilt, Phys. Rev. B 65, 035109 (2001).
+- (63) RESPACK and its user guide can be downloaded from https://sites.google.com/view/kazuma7k6r.
+- Nakamura et al. (2021) K. Nakamura, Y. Yoshimoto, Y. Nomura, T. Tadano, M. Kawamura, T. Kosugi, K. Yoshimi, T. Misawa, and Y. Motoyama, Comput. Phys. Commun. 261, 107781 (2021).
+- Nakamura et al. (2016) K. Nakamura, Y. Nohara, Y. Yosimoto, and Y. Nomura, Phys. Rev. B 93, 085124 (2016).
+- Nakamura et al. (2009) K. Nakamura, Y. Yoshimoto, T. Kosugi, R. Arita, and M. Imada, J. Phys. Soc. Jpn. 78, 083710 (2009).
+- Nakamura et al. (2008) K. Nakamura, R. Arita, and M. Imada, J. Phys. Soc. Jpn. 77, 093711 (2008).
+- Nohara et al. (2009) Y. Nohara, S. Yamamoto, and T. Fujiwara, Phys. Rev. B 79, 195110 (2009).
+- Fujiwara et al. (2003) T. Fujiwara, S. Yamamoto, and Y. Ishii, J. Phys. Soc. Jpn. 72, 777 (2003).
+- Bickers et al. (1989) N. E. Bickers, D. J. Scalapino, and S. R. White, Phys. Rev. Lett. 62, 961 (1989).
+- Bickers and White (1991) N. E. Bickers and S. R. White, Phys. Rev. B 43, 8044 (1991).
+- Sakakibara et al. (2010) H. Sakakibara, H. Usui, K. Kuroki, R. Arita, and H. Aoki, Phys. Rev. Lett. 105, 057003 (2010).
+- Sakakibara et al. (2012) H. Sakakibara, H. Usui, K. Kuroki, R. Arita, and H. Aoki, Phys. Rev. B 85, 064501 (2012).
+- Sakakibara et al. (2014) H. Sakakibara, K. Suzuki, H. Usui, S. Miyao, I. Maruyama, K. Kusakabe, R. Arita, H. Aoki, and K. Kuroki, Phys. Rev. B 89, 224505 (2014).
+- Sakakibara et al. (2013) H. Sakakibara, K. Suzuki, H. Usui, K. Kuroki, R. Arita, D. J. Scalapino, and H. Aoki, J. Phys. Conf. Ser. 454, 012021 (2013).
+- Lichtenstein and Katsnelson (1998) A. I. Lichtenstein and M. I. Katsnelson, Phys. Rev. B 57, 6884 (1998).
+- Yada and Kontani (2005) K. Yada and H. Kontani, J. Phys. Soc. Jpn. 74, 2161 (2005).
+- Mochizuki et al. (2005) M. Mochizuki, Y. Yanase, and M. Ogata, Phys. Rev. Lett. 94, 147005 (2005).
+- Ikeda et al. (2010) H. Ikeda, R. Arita, and J. Kuneš, Phys. Rev. B 81, 054502 (2010).
+- Christiansson et al. (2023) V. Christiansson, F. Petocchi, and P. Werner, Phys. Rev. Lett. 131, 206501 (2023).
+- Sakakibara and Kotani (2019) H. Sakakibara and T. Kotani, Phys. Rev. B 99, 195141 (2019).
+- Sakakibara et al. (2020) H. Sakakibara, H. Usui, K. Suzuki, T. Kotani, H. Aoki, and K. Kuroki, Phys. Rev. Lett. 125, 077003 (2020).
+- Li et al. (2019) D. Li, K. Lee, B. Y. Wang, M. Osada, S. Crossley, H. R. Lee, Y. Cui, Y. Hikita, and H. Y. Hwang, Nature(London) 572, 624 (2019).
+- Dagotto et al. (1992) E. Dagotto, J. Riera, and D. Scalapino, Phys. Rev. B 45, 5744 (1992).
+- Yamazaki et al. (2020) K. Yamazaki, M. Ochi, D. Ogura, K. Kuroki, H. Eisaki, S. Uchida, and H. Aoki, Phys. Rev. Res. 2, 033356 (2020).
+- Kitamine et al. (2020) N. Kitamine, M. Ochi, and K. Kuroki, Phys. Rev. Res. 2, 042032(R) (2020).